# Optimizing Continuous-Wave Pumped Entanglement-based QKD in Noisy Environment

### v2

@@ Abstract
Quantum key distribution (QKD) has emerged as a promising solution to protect current cryptographic systems against the threat of quantum computers. As QKD transitions from laboratories to real-world applications, its implementation under various environmental conditions has become a pressing challenge. Major obstacles to practical QKD implementation are the loss of photons in the transmission media and the presence of extreme noise, which can severely limit long-range transmission. In this paper, we investigate the impact of extreme noise on QKD system parameters, including detector timing uncertainty (jitter), measurement timing shifts, variations in detector deadtime, and rate-dependent detector efficiency. Contrary to manufacturers’ specifications, which assume these parameters to be constant, we demonstrate that these parameters exhibit significant variations in extreme noise conditions. We show that changes in these parameters play a key role in determining system performance in noisy environments. To address these non-idealities, we develop a robust model that can address and adapt to various detector characteristics. In particular, our model is independent of source parameters and can be implemented using data from the detection unit. Our results show that our model is well suited for characterizing and optimizing the performance of the QKD system under noisy conditions.
@@ I INTRODUCTION
Quantum key distribution (QKD) is a method for securely transferring keys between multiple parties using quantum mechanics. QKD uses the no-cloning theory to detect unauthorized access in a two-party cryptographic communication [1, 2] and have been developed in various forms [3, 4, 5, 6, 7, 8], including quantum-entanglement-based QKD protocols[5, 7] and twin-field and decoy-state prepare-and-measure implementations [3, 4]. This system can further be extended to correct instrument-imposed loopholes with the assistance of device-independent QKD protocols [9, 6].
@@
In [20], the authors achieved high performance by optimizing the coincidence window considering detector jitter, photon loss, and source photon’s pair rate. Their experimental study considered a QKD system operating in the telecommunication band and used superconducting nanowire single-photon detectors (SNSPDs). However, non-linear imperfections of the detectors, that can arise due to higher photon detection, have not been considered.
Similarly, the authors in [21] conducted experiments to investigate variations in detector parameters, including dead time, timing jitter, detection timing shift, and afterpulsing effects as a function of detection count rate. They studied these parameters by sending pulsed laser light with well-controlled photon time separation to an actively quenched avalanche photodetector operated in a Geiger mode. Their results showed that these parameters are not constant and demonstrated a method to correct for these variations. This method is useful for mitigating deterioration factors in noisy detection. However, in continuous-wave (CW) sources, photon timing is not well-controlled and is governed by photon statistics.
@@
To characterize and analyze the QKD system under high noise condition, we employed spontaneous parametric down-conversion (SPDC) [22, 23] process as a quantum source and an external broadband noise source features the noise photons. SPDC is a widely used quantum source for QKD realization, and the characteristics of broadband noise resemble noise expected in a practical system. Therefore, our method simulates a realistic scenario. Our findings from a single basis experimental study were extended to a two-basis theoretical model. This model enables the analysis of critical parameters of the QKD, including secure key rate (SKR), quantum bit error rate (QBER), and acceptable noise-photon levels within the coincidence window size of the system. Analyzing these parameters is crucial to improve system performance, particularly in extremely noisy environments.
@@ II.1 Experimental Setup
Fig. 1 shows the experimental setup we use to conduct detector characterization and coincidence measurement analysis. In this experiment, a spatially and spectrally filtered CW laser acts as a pump source with a central wavelength of 405 nm, and bandwidth of 160 MHz, which impinges on a type I configured 6 mm long beta-barium borate (BBO) crystal. As a result, two copolarized and nondegenerate signal and idler photons with wavelengths of 780 and 842 nm are generated, respectively. These two photons are later separated using a dichroic mirror (DM). Signal photons are directly detected by a passively quenched single-photon avalanche photodetector (APD1). The other photon, the idler, passes through a polarizing beam splitter (PBS), and then detected by an actively quenched APD operated in the Geiger mode (APD2). The noise source is directed towards APD2 by sending it from the perpendicular face of the PBS with respect to the idler photons. The noise source we used is a nonpolarized incandescent lamp that demonstrates thermal photon statistics.
In our experiment, only APD2 received noise (APD1 is shielded from noise). Therefore, the measured output from APD2 is a mixture of signal and noise. These two detectors are connected to a timestamp unit (not shown in Fig. 1) to record the time of arrival of individual photons received at both detectors. By analyzing the joint time of arrival of the photons at the detectors, we could obtain a temporal coincidence histogram as shown in Fig.2(a). In the experiment, idler photons travel 1.2 m extra optical path (equivalently, a 4.2 ns shift in the time axis); therefore, the coincidence histogram has an offset from zero point as evident from Fig. 2(a). In the experiment, we used the passive detector made by S15 (with maximum count rate: 120 k cps, timing jitter 250 ns) as APD1, and the Excelitas detector (SPCM-AQRH-10) as APD2. Table 1 shows the important parameters of the Excelitas detector.
@@ II.2 Coincidence Measurement in Noisy Settings
In a typical QKD protocol (e.g., BBM92 protocol [7]), Alice and Bob each require a minimum of two detectors to distinguish orthogonal quantum states. This requirement stems from the protocol’s security mechanism, which relies on the quantum nature of a maximally entangled state and necessitates measurements in two bases. Implementing the protocol requires examining the coincidence counts of measurement events between Alice and Bob using two detector combinations. Therefore, single-basis experimental settings are not sufficient for a complete protocol implementation. However, to study the factors affecting SKR and QBER and their performance changes due to operation in a noisy environment, examining the two detection coincidence events should be sufficient. One key difference is that the photon coincidence rate obtained by a single basis system will be twice that of two basis measurements. Additionally, it is important to note that by considering only one detector at Alice and Bob locations, we are ignoring the polarization error from the basis setting. This is justifiable as this factor will contribute less than of the total errors [24]. However, this leads to a multiplicative change in the overall error rate.
Fig. 2(a) shows a plot generated under various noise conditions (the actual noise values are indicated in the figure). In our analysis, the number of coincidences is determined by accumulating the number of coincidence events every second. By compiling the experimental coincidence values, we can obtain a full coincidence events’ distribution that provides important information about the system. The width of the coincidence events is proportional to the timing jitter of the detector. The Excelitas detector has timing jitter of 0.35 ns according to the manufacturer, which we verify by measuring the full-width at half-maximum (FWHM) of the coincidence distribution in a low noise condition. The S15 detector has slightly less timing jitter, namely 0.25 ns. We observed four distinct changes in the coincidence events as a function of increasing noise.
@@
A proportional time shift in the coincidence events with an increase in noise
@@
An increase of the FWHM of the coincidence distribution as the noise increases
@@
An increase in the floor level of the coincidence events (minimum value of the observed coincidence events) due to noise events
@@
The height of the coincidence histogram started reducing beyond detection threshold of 4.5 Mcps
Not all photon pairs generated in the SPDC process are detected. Due to photon loss during the transmission, the heralding efficiency of the system is degraded and it can be approximated using
@@
where is the number of pairs generated, is the detection efficiency. is inclusive of all loss parameters such as transmission loss, detector efficiency, coupling efficiency. With this, the probability of simultaneously arriving photons in both detectors can be estimated with
@@
Where and are the detection efficiencies of channel and channel , respectively. The detection events reported by the detectors also have noise photons, such as the dark count of the APDs and stray noise entered in the detectors. which can be modeled by modifying Eq. (1) as
@@
where and are the count of noise entered in detector and detector , and are the dark count rates of detector and detector , respectively. In the measurement, there is a random probability that two uncorrelated photon events fall within the same time window, even though they are not from the SPDC process. This is known as accidental coincidence events. These events are the primary source of noise in coincidence-based measurement. Therefore, the rate of accidental coincidence increases with the noise photons. With this, the accidental coincidence is calculated by multiplying single events from two detectors with the chosen coincidence time window which can be written as
where is the chosen coincidence time window. In Fig. 2(a), the accidental coincidence can be estimated by measuring the minimum value read on the coincidence axis. As the dark count of a detector is constant and is known from the manufacturer, the environment-imparted change in the detection events is the only varying parameter. As we do not have another reference for total events and the measured value of detection counts has all parameters, we could estimate accidental coincidences by knowing the coincidence time widow we chose using Eq. (5). With this, we have found that Eq. (5) is consistent with the observed values of and with the chosen .
@@
Typically, real coincidence events are theoretically approximated using a normal distribution of the form [25, 26]
@@
where is the detector jitter and is the measurement time. This method can be applied in numerical simulations to investigate quantum state tomography of entangled qubits with measurements affected by time uncertainty [27]. However, this approximation fails when systems operate in extreme noise regimes. Also, this approximation does not allow one to study individual detector characteristics from the coincidence histogram. In our experiment, where only one detector receives noise-induced clicks, the probabilistic distribution of the coincidence events does not show a clean Gaussian profile in the high noise regime, rendering a normal distribution that does not fit our collected data. Neumann et al. [20] investigated coincidence events using the normal distribution function for various coincidence values, accounting for variations in the Full Width at Half Maximum (FWHM) and shifts caused by dispersion in optical fibers. Using the detector function of the form of Eq. (6), they characterized the changes in the detector parameters using linear functions. In contrast, our observations reveal a polynomial growth in FWHM (see Fig 2b) and a shift in the coincidence time. This discrepancy arises because previous work assumed equal photon pair detection rates across all detectors in a low-noise scenario (in our case, if noise photons entered <4.5 Mcps, the widening of FWHM will appear almost linear). Consequently, in their study, FWHM changes and coincidence peak shifts appeared linear.
@@
Our results show noise photon can be responsible for shift in the detection timing, particularly its evident when detection rates are imbalanced between detectors. To address this, we should be able to cater to the changes in individual detectors reflect in the theoretical fit. By incorporating noise-induced variation parameter with exponential distribution coincidence histogram, we could detect changes in the individual detectors as follows
@@
where is a coefficient used to accommodate the change in the width of coincidence events, is the dead time of the detector and is the centroid value of the coincidence histogram. will provide the FWHM in the time domain. By convoluting detector function in Eq. (6) with and , we can obtain a complete coincidence histogram function of the form:
@@
where and are the timing uncertainty (jitter) of the detectors used at signal and idler arms, respectively, is the shift of centroid point of coincidence histogram from the origin, and “” represents the complimentary error function. Detector timing jitter is predominantly responsible for the width of the coincidence events. The timing uncertainty of the detectors (jitter) is responsible for the width of the coincidence histogram plots. If the temporal coherence of the collected photons is significantly greater than the jitter of the detection system, the measured spectral profile will reflect the true spectral characteristics of the photons more closely [28, 29]. The coherence time of our photon pairs is about 21.5 fs (corresponding to the spectral bandwidth of 30 nm), which is negligible and overshadowed by the timing jitter of the detectors. Therefore, timing uncertainty due to this factor could be omitted. Additionally, the temporal resolution of the timestamp unit also contributes to the width of coincidence events if it is greater than the timing jitter. Since the temporal resolution of our timestamp unit is 81 ps, we can omit this factor as well. When the noise is low, we can approximate . As noise increases, we adjusted the value of to fit with the experimental data. Theoretical fit for the collected experimental data is shown in Fig. 2(a). The change in the FWHM in different noise settings is shown in Fig. 2(b). We observe that is proportional to the FWHM of the coincidence histogram. The experimentally obtained shift of the start-stop histogram was directly incorporated into Eq. (9) during data fitting. One parameter that largely depends on the shape of the coincidence histogram is the variation in the dead time of the detectors due to high photon detection events. We have investigated the change in dead time with increased event rates in the detector. We used experimentally obtained values (discussed in Section II.3) in our model.
| Parameter | SPCM-AQRH-10 |
| deadtime | 22 ns |

### v4

# Optimizing Continuous-Wave-Pumped Entanglement-based QKD in Noisy Environments

@@ Abstract
Quantum key distribution (QKD) has emerged as a promising solution to protect current cryptographic systems against the threat of quantum computers. As QKD transitions from laboratories to real-world applications, its implementation under various environmental conditions has become a pressing challenge. Major obstacles to practical QKD implementation are the loss of photons in the transmission media and the presence of extreme noise, which can severely limit long-range transmission. In this paper, we investigate the impact of extreme noise on QKD system parameters, including timing jitter, rate-dependent timing shifts, changes in effective detector dead time, and rate-dependent detection efficiency. Contrary to manufacturers’ specifications, which assume these parameters to be constant, we demonstrate that these parameters exhibit significant variations in extreme noise conditions. We show that changes in these parameters play a key role in determining system performance in noisy environments. To address these nonidealities, we develop a model that adapts to detector-dependent timing distortions and recovery effects. In particular, our model is independent of source parameters and can be implemented using data from the detection unit. Our results show that the model enables reliable characterization and optimization of QKD performance under strong noise.
@@ I INTRODUCTION
Quantum key distribution (QKD) is a method for securely transferring keys between multiple parties using quantum mechanics. QKD exploits quantum principles (including the no-cloning theorem) to detect eavesdropping in two-party cryptographic communication [1, 2] and has been developed in various forms [3, 4, 5, 6, 7, 8], including quantum-entanglement-based protocols [5, 7], decoy-state prepare-and-measure schemes [3, 4], and more recently twin-field QKD protocols [9, 8]. This system can further be extended to correct instrument-imposed loopholes with the assistance of device-independent QKD protocols [10, 6].
@@
In Ref. [21], authors achieved high performance by optimizing the coincidence window considering detector jitter, photon loss, and source photon’s pair rate. Their experimental study considered a QKD system operating in the telecommunication band and used superconducting nanowire single-photon detectors (SNSPDs). However, they did not consider nonlinear detector imperfections that can arise at high detection rates. In addition, we observe a pronounced shift in the coincidence timing in the extreme-noise regime. This shift is rooted in photon arrival statistics and in rate-dependent recovery dynamics of silicon-based avalanche photodetectors.
Similarly, the authors in [22] conducted experiments to investigate variations in detector parameters, including dead time, timing jitter, detection timing shift, and afterpulsing effects as a function of detection count rate. They studied these parameters by sending pulsed laser light with well-controlled photon time separation to an actively quenched avalanche photodetector operated in a Geiger mode. Their results showed that these parameters are not constant and demonstrated a method to correct for these variations. This method is useful for mitigating deterioration factors in noisy detection. However, this approach is not applicable to continuous-wave (CW) pumped QKD systems, where photon arrival times are not externally controlled but governed by photon statistics.
@@
To characterize and analyze the QKD system under high noise conditions, we used the spontaneous parametric down-conversion (SPDC) [23, 24] process as a quantum source and an external broadband noise source features the noise photons. SPDC is a widely used quantum source for QKD realization, and the characteristics of broadband noise resemble noise expected in a practical system. Therefore, our method simulates a realistic scenario. Our findings from a single basis experimental study were extended to a two-basis theoretical model. This model enables the analysis of critical parameters of the QKD, including secure key rate (SKR), quantum bit error rate (QBER), and acceptable noise-photon levels within the coincidence window size of the system. Analyzing these parameters is crucial to improve system performance, particularly in extremely noisy environments.
@@ II.1 Experimental Setup
Our experimental setup is shown in Fig. 1. We employ a single-frequency CW laser (OBIS, Coherent) operating at a central wavelength of 405 nm as the pump laser. The pump beam is spectrally cleaned using a bandpass filter with a bandwidth of 10 nm, followed by a variable neutral-density (ND) filter to control the pump power. The beam is then coupled into a single-mode optical fiber to ensure a well-defined spatial mode.
At the fiber output, the pump beam is focused into a type-0 periodically poled potassium titanyl phosphate (PPKTP) crystal using a plano-convex lens with a focal length of 175 mm (L175 in Fig. 1). A half-wave plate is placed immediately after the fiber exit to optimize the pump polarization for efficient type-0 quasi-phase-matched SPDC. The crystal is aligned for collinear generation of photon pairs near the degenerate wavelength of 810 nm by tuning the temperature of the oven to Celsius.
After the PPKTP crystal, residual pump light is removed using a long-pass filter with a cutoff wavelength of 650 nm. The emerging collinear SPDC field, which exhibits a conical emission profile, is subsequently collimated using a lens with a focal length of 75 mm (L75 in Fig. 1). A narrowband interference filter centered at 810 nm is then used to further reject any remaining pump photons and background light.
The collinear SPDC beam is spatially divided into two symmetric paths using a knife-edge prism positioned to bisect the transverse intensity profile of the SPDC beam, thereby separating the spatial mode into two channels. In both channels, photons are coupled into single-mode fibers and detected using identical fiber-coupled avalanche photo detectors (APDs, Excelitas SPCM-AQRH-14-FC). The Salient features of this detector are summarized in Table 1).
The two detectors collect The photon arrival times are recorded using a Swabian Instruments Time Tagger Ultra (timing resolution 8 ps), but histogram bin width we used is 85 ps. With this arrangement, The noise-free coincidence rate of the channel-1 and channel-2 is calculated to 6321 counts per second (cps) when the measurement window selected to be 16.915 ns, and its heralding efficiency is 3.2 throughout the coincidence measurement. In Channel 2, controlled noise is deliberately introduced by mixing the SPDC channel-2 photons with light from a broadband source (Thorlabs SLS201L/M). This noise source is a Tungsten-Halogen Light Source. Since this source exhibits thermal light characteristics with broad spectral width, and there can simulate sunlight. A polarizing beam splitter is used to efficiently mix the noise and channel-2 signal. Stack of neutral-density filters are placed in the noise arm that allows well control the noise intensity. This configuration enables systematic investigation of detector response under varying noise conditions while maintaining identical detection hardware in both channels.
In our experiment, only APD at channel-2 received noise (channel-1 is shielded from noise). Therefore, the measured output from channel-2 is a mixture of signal and noise. These two detectors are connected to a time tagger to record the time of arrival of individual photons received at both detectors. By analyzing the joint time of arrivals of the photons at the detectors, we could obtain a temporal coincidence histogram as shown in Fig.2(a) (the python code to achieve the coincidence measurement is given in the appendix. A). This temporal coincidence measurement is a basis for most of the discrete variable QKD experiments. In our setup experiment, idler photons travel m extra optical path (equivalently, a ns in time domain); therefore, the coincidence histogram has an offset from zero point as evident from Fig. 2(a).
detector
@@ II.2 Coincidence Measurement in Noisy Settings
In a typical QKD protocol (e.g., BBM92 protocol [7]), Alice and Bob each require a minimum of two detectors to distinguish orthogonal quantum states. This requirement stems from the protocol’s security mechanism, which relies on the quantum nature of a maximally entangled state and necessitates measurements in two bases. Implementing the protocol requires examining the coincidence counts of measurement events between Alice and Bob using two detector combinations. Therefore, single-basis experimental settings are not sufficient for a complete protocol implementation. However, to study the factors affecting SKR and QBER and their performance changes due to operation in a noisy environment, analyzing a single coincidence channel is sufficient to isolate the dominant timing and noise effects. One key difference is that the photon coincidence rate obtained by a single basis system will be twice that of two basis measurements. Additionally, it is important to note that by considering only one detector at Alice and Bob locations, we are ignoring the polarization error from the basis setting. This is justifiable as this factor will contribute less than of the total errors [25]. Additionally, this only leads to a multiplicative change in the overall error rate.
Fig. 2(a) shows a plot generated under various noise conditions (the actual noise values are indicated in the figure). In our analysis, the number of coincidences is determined by accumulating the number of coincidence counts in every second. By compiling the experimental coincidence values, we can obtain a full coincidence counts’ distribution that provides important information about the system operating in a noise setup. For APD-based detection, the coincidence-peak width is primarily determined by detector timing jitter (convolved across both detectors). Our detectors have timing jitter of 350 ps (according to the manufacturer), which we verify by measuring the full-width at half-maximum (FWHM) of the coincidence distribution in a low noise condition. We observed four distinct changes in the coincidence counts as a function of increasing detection counts (driven by noise in this case).
@@
A systematic shift of the coincidence peak, which becomes evident above 1 Mcps and increases with count rate.
@@
Broadening of the coincidence peak (FWHM increase), also evident above 1 Mcps.
@@
An increase in the coincidence baseline due to accidental coincidences.
@@
A reduction in peak amplitude at high count rates, consistent with rate-dependent efficiency degradation.
The coincidence peak shift originates from rate-dependent distortion of the single-detector timing response. As noise is injected into one detection channel, the inter-arrival time distribution becomes increasingly truncated by detector dead time, leading to a delayed and asymmetric temporal response. Since the coincidence histogram represents a convolution of the two detector response functions, the asymmetric distortion of one channel shifts the location of maximum temporal overlap, resulting in a systematic displacement of the coincidence peak. If we account loss of photon during the transmission, we can estimate the photon detection rate on the detector using
@@
where is the number of pairs generated in the crystal, is the detection efficiency. We assume transmission loss, and other loss parameters negligible. With this, coincidence detection probability of photons at channel-1 and channel-2 can be estimated following
@@
Where and are the quantum efficiencies of detectors at channel-1 and Channel-2, respectively. The detection events registered by the detectors also have noise counts, such as the dark count of the APDs () and stray photons entered in the detectors (). which can be modeled by modifying Eq. (1) into
@@
where and are the count of noise entered in channel-1 and channel-2, and are the dark count rates of detector- and detector-, respectively. In the measurement, there is a random probability that two uncorrelated photon events fall within the same measurement time window. This is known as accidental coincidence counts. These events are the primary source of noise in coincidence-based measurements. The rate of accidental coincidence increases with the noise photons following the relation
where is the selected coincidence measurement time window. In experiments, the accidental coincidence can be estimated by measuring y-axis offset of start-stop histogram (see Fig. 2(a)). Unlike genuine coincidences, accidental coincidences produce a nearly time-independent background in the start–stop histogram. As the dark count of a detector is constant and is known from the manufacturer, the environment-impacted change in the detection events is the only varying parameter.
@@
Typically, the coincidence measurements are theoretically approximated using a normal distribution of the form [26, 27]
@@
where is the detector timing jitter and is the measurement time. This method can be applied in numerical simulations to investigate quantum state tomography of entangled qubits with measurements affected by time uncertainty [28]. However, this approximation fails when systems operate in high noise regimes. Also, this approximation does not allow one to study individual detector characteristics from the coincidence histogram. In our experiment, where only one detector receives noise-induced clicks, the probabilistic distribution of the coincidence counts does not show a clean Gaussian profile in the high noise regime, rendering a normal distribution that does not fit our collected data. Neumann et al. [21] investigated coincidence counts using the normal distribution function for various coincidence values, accounting for variations in the Full Width at Half Maximum (FWHM) and shifts caused by dispersion in optical fibers. Using the detector function of the form of Eq. (6), they characterized the changes in the detector parameters using linear functions. In contrast, our observations reveal a polynomial growth in FWHM (see Fig 2b) and a shift in the coincidence time. This discrepancy arises because previous work assumed equal photon pair detection rates across all detectors in a low-noise scenario. Consequently, in their study, FWHM changes and coincidence peak shifts appeared linear.
@@
Our results show external photons can induce the shift in the detection timing. In particular, this becomes evident when detection rates are imbalanced. To efficiently address this, it is therefore necessary to introduce a model that could cater changes in the individual detectors in the coincidence profile. By incorporating noise-induced variation parameter, one could detect changes in the individual detectors using the formulation of
@@
where is a coefficient used to accommodate the change in the width of coincidence counts, is the dead time of the detector and is the centroid value of the coincidence histogram. will provide the FWHM in the time domain. By incorporating detector function in Eq. (6) with and , we can obtain a complete coincidence histogram function of the form:
@@
where and are the timing uncertainty (jitter) of the detectors used at channel-1 and channel-2, respectively, is the shift of centroid point of coincidence histogram from the origin, and “” represents the complementary error function. Detector timing jitter is predominantly responsible for the width of the coincidence counts. If the temporal coherence of the collected photons is significantly greater than the jitter of the detection system, the measured spectral profile will reflect the true spectral characteristics of the photons more closely [29, 30]. The coherence time of our photon pairs is about 60 fs (corresponding to the spectral bandwidth of approximately 16 nm, this value is taken from Ref. [31]), which is negligible and overshadowed by the timing jitter of the detectors. Therefore, timing uncertainty due to spectral width of the SPDC photons could be omitted. Additionally, the temporal resolution of the time tagger also contributes to the width of coincidence counts if it is greater than the timing jitter. Since the temporal resolution of our time tagger is 8 ps, since this is much less than , we can omit this factor as well. In a noise-free operation of the detectors, we can approximate . As noise increases, we adjusted the value of to fit with the experimental data.
II.3 Detection timing shift and FWHM widening in start-stop histogram
Since we saturate only one detector, the coincidence peak measures the mean relative delay between channel-1 and channel-2. At low count rates, each avalanche is fully quenched and the detector fully recovers between detection events. The discriminator threshold is crossed at a fixed point in the avalanche buildup. In this case, the detection latency is deterministic with fixed jitter. Mathematically, detection time is:
where is the fixed latency, and is the stationary jitter (Gaussian). As a result in coincidence, we observe stable peak position, constant width (constant FWHM) of start-stop coincidence measure, maximum peak height, and low accidental background. When we increase the noise rate in detector-2, this detector rarely fully recovers. Many photons arrive during the recovery tail. Consequently, effective bias voltage is lower than nominal. Avalanche growth is slower. Therefore, detection latency becomes state-dependent. The detection delay now depends on the time since the last avalanche: . This function increases as recovery worsens, and saturates at high rate. In our case detector-1 is stable and detector-2 develops a positive mean latency shift in a count regime:
Thus the rate-dependent delay we observe is:
Because recovery is related to Poisson statistics, the probability that at least one photon appears in the time "t" is given by
where R is the photon rate. Due to characteristic recovery time of the detector, it cannot detect photons if , where is the characteristic recovery time of the detector. Thus, Eq. 14 modify into
The coincidence peak shift we observed should follow this behavior with intrinsic delay incorporated in it. Thus, the shift we observed can be estimated using
where "A" is the amplitude of this effect, is the characteristic rate. We also observed widening of start-stop histogram. This is because not all detections are delayed by the same amount. The recovery state varies randomly event-to-event, so the detection latency distribution broadens. Thus, detector-2 has larger variance in detection time. In low count rate measurement the Coincidence width obeys
Since only increases, FWHM increases, and peak becomes asymmetric. This extra jitter saturates at high rate. Since both peak shift and increase in FWHM follow same recovery dynamics, we expect both follows the trend. Following this, we can observe widening of the FWHM of the start-stop coincidence profile by following
The experimentally extracted FWHM and peak shift are shown in Fig. 2(b). Both are well fitted by a saturating recovery model with a shared characteristic rate, supporting the hypothesis of a common recovery-driven mechanism. This summary, injecting noise photons into one detector introduces a rate-dependent modification of the detector recovery dynamics. At high count rates, photons are detected before full recovery, leading to both an increased mean detection latency and enhanced variability in the detection time. In coincidence measurements, this manifests as a shift of the coincidence peak, a reduction in peak amplitude due to reduced effective detection efficiency, an increase in the accidental background, and a broadening of the coincidence peak. Both the peak shift and the FWHM exhibit a saturating dependence on the noise count rate with a shared characteristic rate, indicating a common physical origin in detector recovery effects.
| Parameter | SPCM-AQRH-14-FC |
| Dark count | 39 cps |
| maximum count | 38.1 MHz |
| dead time | 23 ns (@ < 1MHz) |